OpenStudy (anonymous):

Students are given a list of ten vocabulary words to learn. In how many ways could four of the words be listed on a test? Is this permutation or combination??

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

It'd be worth looking up "permutation" and "combination" to get a better idea of what each means. I'd call the present situation a "combination:" the number of ways in which we could choose from 10 distinct objects 4 at a time. Here, n = 10 (n is the number of distinct objects we have from which to choose) x = 4 (the number of objects taken at any one time) C We want to evaluate 10 4. Have you any previous experience in evaluating combinations such as this?

OpenStudy (mathmale):

\[_{10}C _{4}=\frac{ 10! }{ 4!(10-4)! }\]

OpenStudy (mathmale):

Note: n! means "n factorial," and is equal to n(n-1)(n-2)(n-3)....1. So: 4! = 4*3*2*1=24

OpenStudy (anonymous):

thanks so much! I knew how to do both combinations and permutations I just could not figure out which this one was to do it!